The Pettit Corporation has annual credit sales of $2 million. Current expenses for the collection department are $30,000, bad debt losses are 2 percent, and the DSO is 30 days. Pettit is considering easing its collection efforts so that collection expenses will be reduced to $22,000 per year. The change is expected to increase bad debt losses to 3 percent and to increase the DSO to 45 days. In addition, sales are expected to increase to $2.2 million per year. Should Pettit relax collection efforts if the opportunity cost of funds is 12 percent, the variable cost ratio is 75 percent, and its marginal tax rate is 40 percent? Assume 365 days per year.

. Use the income statement approach to find the incremental bad debt losses

. Use the income statement approach to find the incremental cost of carrying receivables

. Use the income statement approach to find the incremental pre-tax profits from this proposal

. Use the incremental analysis approach to find the incremental bad debt losses

. Use the incremental analysis approach to find the incremental cost of carrying receivables

. Use the incremental analysis approach to find the incremental pre-tax profits from this proposal
Answer #1
Present (A) Proposed (B) Incremental (A)-(B)
current sales 2000000 New sales 2200000 200000
credit period 30 days credit period 45 days
Receivables 164384 (sales*30/365) Receivables(sales*45/365) 271233 106849
bad debt-2% 40000 bad debt-3% 66000 26000
carrying cost of receivables-12%*164384 19726 carrying cost of receivables=12%*271233 32548 12822
Contribution 500000 Contribution 550000 50000
Variable cost 75%
carrying cost 12%
Tax rate 40%
Increase in Sales 200000
Increase in contribution 50000 (contribution=100-75=25)*inc in sales)
Solution:
a Incremental bad debts
26000 (as per above working)
b Incremental cost of carrying debts
12822 (as per above working)
c Incremental Pretax profits:
Increase in Contribution 50000 (as per working note)
(increase in profits)
Incremental contribution is equal to pretax profits)
Note:
In both Income statement approach (means estimating the amount based on ratios/judgements without referring to actual balance in accounts)
and incremental analysis approach, the solution would be same while calculating the incremental bad debts, pretax profits etc., because in both,
incremental figures are considered